find the value of m if (-2,0) and (4,m) are coordinates of twp points on a straight line whoae slope is 5/6
Answer by mananth(16946)

(-2,0) and (4,m) are coordinates of twp points on a straight line whoae slope is 5/6
x1 y1 x2 y2
-2 0 4 m

slope m = (y2-y1)/(x2-x1)
( m -0 )/( 4 - -2 )
( m / 6 )
m/6 = 5/6
m= 5*6/6
m=5